OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (−1, 6) and has a slope of −3.

OpenStudy (anonymous):

ill give you an example, and see if you get it:

OpenStudy (anonymous):

alright thank you.

OpenStudy (anonymous):

* Find the equation of the straight line that has slope m = 4 and passes through the point (–1, –6). They've given me m = 4, x1 = –1, and y1 = –6. I'll plug these values into the point-slope form, and solve for "y=": y – y1 = m(x – x1) y – (–6) = (4)(x – (–1)) y + 6 = 4(x + 1) y + 6 = 4x + 4 y = 4x + 4 – 6 y = 4x – 2 Cop

OpenStudy (anonymous):

this is called the point slope formula y – y1 = m(x – x1)

OpenStudy (anonymous):

please add as fan :DD

OpenStudy (anonymous):

im going to try to figure it out right now. please tell me if i get the right answer.

OpenStudy (anonymous):

yes, i will

OpenStudy (anonymous):

oh, ignore the "cop" in the end. that not part of it

OpenStudy (anonymous):

y-6=-3(x-(-1) y+6=-3(x+1) y+6=3x+1 y=3x+4+1 y=3x+5

OpenStudy (anonymous):

wait is it y=-3x+3?

OpenStudy (anonymous):

ok, so you knew what you were doing, but you messed up on the second line. you dont make -6, to +6. They did that, because it was already a negative, so it becomes a positive( y-(-3) become y+3

OpenStudy (anonymous):

your second answer is right

OpenStudy (anonymous):

okay thank you! :) can you help me with giving me some more examples to other problems that i give you?

OpenStudy (anonymous):

sure, but i gtg , in a sec

OpenStudy (anonymous):

alright, just tell me when your back on.

OpenStudy (anonymous):

heres a great math website that gives you problems, and you can see your the answer after your finished solving it

OpenStudy (anonymous):

practice alot of these, cuz practice makes perfect :DD

OpenStudy (anonymous):

thank you. :)